OpenStudy (anonymous):

The volume V of a tumor that commonly develops in canines is directly proportional to the cube of the time t that the tumor has been growing. If a 3-month old tumor has a volume of 1.89 cm3, what will be the size of a tumor that is 4 months old?

OpenStudy (thewayiam983):

Vt^3=some constant k (1.89)(3)^3=k = 51.03 51.03=Vt^3 = V(64) V=51.03/64=.797 cm^3

OpenStudy (anonymous):

options are 4.48cm^3, 3.48cm^3, 4.84cm^3 or 4.40cm^3

OpenStudy (thewayiam983):

oooh i may have done that backwards. directly proportional. V/t^3=k 1.89/27=k=.07 .07=V/64 V=.07*64=4.48

OpenStudy (thewayiam983):

my apologies

OpenStudy (anonymous):

np tyvm